ch3-04 student_name.xis (h Print each completed worksheet in Vallle printed in the lower left footer and the file name in the lower right foote Chapter 3 Case Problem 2: WINE DEPOT You are to create Wine Depots financial analyses as of December 31, 2020, and as of December 31, 2021. Following the Chapter 3 examples, use the stu- dent file ch3-05.xls to create a vertical and horizontal analysis of the balance sheet and income statement in the columns provided. (Note: The horizontal analysis will compare December 31, 2020 with December 31, 2021.) Also cre- ate a pie chart of expenses for the year ended December 31, 2021; a column chart of expenses for the years ended December 31, 2020, and December 31,

2021; and a ratio analysis as of December 31, 2021. Save the file as ch3-05_student name.xls (replacing student_name with your name). Note: You will have to use Excels help feature to create the column chart, because the columns are not adjacent to one another as in the chapter example. Use what- ever chart layout you prefer. Print each completed worksheet in Value view, with landscape orientation, scaling to fit to 1 page wide by 1 page tall, and with your name and date printed in the lower left footer and the file name in the lower right footer. Use Excels help feature to move charts on your worksheet so that they print on a separate page.

Number Font Al XV Wine Depot, Inc G HI Wine Depot Inc. Income Statennt for the period ended December-2020 %Sales December-2021 %Sales %Change 6 Sales 7 Less: Cost of Goods Sold 8 Gross Margin 9 Expenses: 10 Advertising 11 Marketing 12 Depreciation 13 Interest 14 Salary 15 Wages s 375,000 150,000 225,000 S 425,000 191,250 S 8,000 6,000 9,000 3,000 52,000 18,000 1,200 s 18,250 3,000 8,000 1,500 85,000 15,000 7,000 Supplies 17 Utilities 18 Net income before taxes 19 Income taxes 0 Net income 2,800 100,000 125,000 37,500 S 87,500 8.500146,250 87,500 26,250 S 61,250 28 30 32 Income Statement alance Sheet Ratio Analysis O Type here to search 0峊eoa Φ @粗v Wine Depot, Inc. Wine Depot, Inc Balance Stheet as of December 2021 %Assets %Change December-2020 %Assets Assets Current Assets: S 7,000 3,500 8,500 10,500 500 S 5,000 7 Cash Short-Term Investments 6,000 9 Accounts Receivable 6,500 1,800 22,000 85,000 10 Inventory 11 Prepaid Expenses 12 Supplies 13 2.300 32.300 1 03,700 14 Equipment 15 Less: Accumulated Depreciation 8,000 77000 16,000 87.700 $120,000 16 Total 17 18 Liabilities and Stockholders Equity 9 Liabilities 20 Current Liabilities 21 Accounts Payable 22 Income Tax Payable 3 Notes Payable Equipment 24 Notes Payable- Other S 3,000 1,000 S 16,000 9,000 18,000 3,000 3,75010,750 Payroll Liabilities 2,00051,000 Stockholders Equity 27 Common Stock 28 Retained Earnings 29 Total 1,000 1,000 47.000 48,000 $ 99,000 10.250 109.250 $ 120,000 31 Income Statement Balance Sheet Ratio Analysis O Type here to search

b Copy Format Painter Number Alignment Clipboard A4 Wine Depot, Inc. Ratio Analysis December 2021 5 Profitability 6 Return on owners investment Return on total investment 8 Profit margin 9 Gross margin 10 Liquidity 11 Current ratio 12 Quick ratio 13 Receivable tumover 14 Inventory tumover 15 Solvency 16 Debt-to-equity 17 Liability 18 23 28 30 31 e Statement Balance Sheet Ratio Analysis O 0 Type here to search
Answer #1

wine depot inc.

Income Statement

for the period ended

december 2020 %sale december 2021 %sale %change
sales $375000 100 425000 100 13.33
less: cost of goods sold 150000 40 191250 45 27.5
gross margin 225000 60 233750 55 3.88
expenses
advertising $8000 2.133 $18250 4.29 128.125
marketing 6000 1.6 3000 .71 -50
deperication 9000 2.4 8000 1.88 -11.11
interest 3000 0.8 1500 .35 -50
salary 52000 13.87 85000 20 63.46
wages 18000 4.8 15000 3.53 16.667
supplies 1200 0.32 7000 1.65 483.333
utilities 2800 100000(26.67) .75 8500 146250(34.41) 2 202.57
net income before taxes 125000 33.33 87500 20.59 -30
income tax 37500 10 26250 6.18 -30
net income $87500 23.33 61250 14.41 -30

wine depot

balance sheet as of

assets december 2020 %assets december 2021 %assets %change
current assets
cash $5000 5.05 7000 5.833 40
short term investment 2000 2.02 3500 2.92 75
account receivable 6000 6.06 8500 7.08 41.67
inventory 6500 6.57 10500 8.75 61.54
prepaid expenses 700 0.71 500 .42 -28.57
suuplies 1800 22000 1.82 2300 32300 1.92 27.78
Equipments 85000 103700
Less deperication 8000 77000 77.78 16000 87700 73.08 13.9
Total 99000 100 120000 100 21.21
Liability and stock holders equilty
liabilities
current liabilities
accounts payable $16000 16.16 $3000 2.5 -81.25
income tax payable 6000 6.06 1000 0.833 -83.33
notes payable - equipments 9000 9.09 - 0 -100
notes payable - others 18000 18.18 3000 2.5 -83.33
payroll liabilities 2000 51000 2.02 3750 10750 3.125 87.5
stockholders equilty
common stock 1000 1.01 1000 .83 0
retained earning 47000 48000 47.47 108250 109250 90.21 130.32
total $99000 100 $120000 100 21.21

wine depot inc.

ratio analysis

december 2021

Profitability ratio

return on owners investment = net income/shareholders equity = 61250/109250 = 0.56

return on total investment = return before interest and tax / total assets = 89000/120000 = 0.74

profit margin = net income/net sale = 61250/425000 = 0.144

gross margin = gross profit/ net sale = 233750/425000 = 0.55

liquidity ratio

current ratio = current assets/current liability = 32300/10750 = 3.00465

quick ratio = current assets - stock / current liability = 19000/10750 = 1.77

average receivable = opening receivable + closing receivable / 2 = 6000+8500/2 = 7250

receivable turnover ratio = netsale/ average receivable = 425000/7250 = 58.62

inventory turnover= cogs/average stock = 191250/8500 = 22.5

average stock = opening stock + closing stock / 2 = 6500+10500/2 = 8500

solvency ratio

debt to equity = total liability / total share holder equity = 10750/109250 = 0.098

liability ratio = total liability / total assets = 10750/120000 = 0.0896